When saving = Investment, show that the trade balance equals the government spending balance?


1
Expert's answer
2022-02-07T16:09:24-0500

"M-X=I-S-G"

Since"I=S"

"I-S=0"

Therefore"M-X=-G"

Dividing by -1 on both sides we are left with:"X-M=G"